A circular coil has a 8.24 cm radius and consists of 39.0 closely wound turns of wire. An externally produced magnetic field of magnitude 3.08 mT is perpendicular to the coil. (a) If no current is in the coil, what magnetic flux links its turns? (b) When the current in the coil is 5.53 A in a certain direction, the net flux through the coil is found to vanish. What is the inductance of the coil? (a) Number Enter your answer for part (a) in accordance to the question statement Units Choose the answer for part (a) from the menu in accordance to the question statement (b) Number Enter your answer for part (b) in accordance to the question statement Units Choose the answer for part (b) from the menu in accordance to the question statement
Answer #1

Answer Explanation:

    (a)   ΦB = NBA
                = (39)(3.08mT)(3.14*(0.0824m)2) (since A =πr2)
                = 2.5609 mWb

    (b)   L =ΦB /i
              =(2.5609 mWb)/(5.53A)
              = 0.4631 mH

-------------------------------------------------------------------------------------------------

Hope this helps!

Please Upvote life saver...

Thank you.